answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Which is more consistency arthematice mean is 110 and standard deviation is 25 and arthematic mean is 90 and standard deviation is 15?

The standard deviation is a number that tells you how scattered the data are centered about the arithmetic mean. The mean tells you nothing about the consistency of the data. The lower standard deviation dataset is less scattered and can be regarded as more consistent.

How do you compute the probability of an event?

There are two main ways:

One is to calculate the theoretical probability. You will need to develop a model for the experiment and then use the laws of science and mathematics to determine the probability of the event (subject to the model's assumptions).

A major alternative is the empirical or experimental method. This requires performing the trial many times. The probability of the event is estimated by the proportion of the total number of trials which result in the outcome of interest occurring.

Can you downgrade your planet fitness black card?

yes you can all you have to do is pay the startup fee for a white card which is normally ten dollars.

How would you calculate the probability you'll have experienced a rainy day at some point in the last seven days if there's been a 10 percent chance of rain each day?

Rainy days are not independent events, but for theis question you will need to assume that they are. Otherwise, there is insufficient information to answer.

Prob(a day being rainy) = 0.1

so Prob(a day being not rainy) = 1 - Prob(a day being rainy) = 1 - 0.1 = 0.9

Therefore, Prob(7 days being no rainy) = (0.9)7

and so, Prob(At least one rainy day in 7) = 1 - Prob(no rainy days in 7)

= 1 - (0.9)7 = 0.5217 (approx).

What is the probable function of your wisdom teeth?

Long ago it was thought that the probable function of your wisdom teeth was to enable the chewing of rough food. However, it would appear that evolution has rendered the wisdom teeth to be pointless and no longer serves a purpose.

What is the probability of getting at least 2 tails in 4 tosses of a fair coin?

The probablility of getting 2 tails in 4 tosses of a fair coin is most likely 50%, 2/4=1/2, or .50.

How do you calculate standard deviation?

standard deviation is the positive square root of mean of the deviations from an arithmatic mean X denoted as sigma.

sigma=sqrt {(sum(x-X)^2)/n}

Probability of a tsunami on the east coast of the US?

Very close to 1.

Over the next 4.5 billion years that the earth is expected to remain in existence, such an event is almost certain.

How do you estimate 974?

no need for estimates i'll give u the answer . the answer is 1,000

Why would an IQ test be developed so most people would score between 85 and 115?

An IQ test is simply a (somewhat flawed) means of assessing a person's "relative intelligence". It is important that IQ is not extremely important when working with such a measure in psychology.

Now, the IQ test is designed such that scores follow a pattern known as a "normal distribution". This is not simply an expected distribution! This distribution is very important in the study of statistics, and has many applications to scientific disciplines (such as psychology).

A standard deviation is essentially a measure of distance from the mean. For example, say that the average person lives to be 75, with a standard deviation of 4. This means that a person who lives to be 79 would be one standard deviation above the mean, and a person who lives to be 67 would be two standard deviations below the mean.

As a table depicting the normal distribution can show, one standard deviation above the mean is at the 84th percentile, while one standard deviation below the mean is at the 16th percentile (two stds above: 97.5th, two stds below: 2.5th)

This means that an IQ score of 115 is at the 84th percentile, while an IQ score of 85 is at the 16th percentile. 68% of all people are within this range, and that is "most" people.

What are 3 different names for 20?

1. a score

2. twenty, vingt, zwanzig, etc

3. vicenary (adjective), meaning consisting of, or relating to, or based on 20.

e.g. "The constitution requires that such serious matters must be considered by a plenary session of the house of representative, whereas lesser matters such as local by-laws can be decided by a vicenary committee, which must include the president, vice-president, and the appointed ministers of state representing the various interests of the state and its citizens."

How many ways can four cars be parked in a row of four parking spaces?

There are 16 ways that four cars can be parked in a row of four parking spaces. You would multiply the number of cars by the number of spaces.

A coin is tossed If it comes out heads a die is then tossed once if a tail appears the coin is tossed one more time List the possibilities as to how many outcomes are possible?

The answer in the back of the book says that it is: TT, TH, H1, H2 H3, H4, H5, H6= 8 possibilities. But I don't understand what that means or how the answer was found. Help?

My dad has vitiligo and he got it when he was in his 40's Im 30 what are the chances I will get this also?

Its not necessary that if parents have vitiligo then their children will definitely have vitiligo. But there are chances of its inheritance. Some environmental factors may work as triggering agents for vitiligo.

What methods would you use to determine the probability of rolling a 3 or a 5 on each of the three consecutive trials with a standard die?

If you want to ask questions about the "following", then I suggest that you make sure that there is something that is following.

When you toss a coin what is the probability it will land on heads?

the probability is actually not quite even. It would actually land heads 495 out of 1000 times because the heads side is slightly heavier

How many different ways can 50 players in a marching band be arranged in rectangle arrangements?

Well, let's see. There has to be a whole-number of players in each row and

each column (rank and file), so the dimensions ofthe rectangle have to be

the whole-number factors of 50.

So the only arrangements we see, without leaving any empty places, are

1 x 50

2 x 25

5 x 10

players.